# Relative monodromy of abelian logarithms for finite covers of universal families

### v2

@@ 2.4 Relative monodromy of abelian logarithms
Let’s continue considering an abelian scheme $\mathcal{A}\rightarrow S$ and a rational non-zero section $\sigma:S\rightarrow\mathcal{A}$. Generally neither a basis of periods nor a logarithm $\log_{\sigma}$ can be globally defined on the whole of $S(\mathbb{C})$ (see 2.7 and 2.10 below), but obviously they can be globally defined on the universal cover of $S$. Studying the related monodromy problems corresponds to finding out the minimal (unramified) cover of $S$ on which both a basis of the periods and a logarithm of the section can be defined. With this in mind, we first call $S^{*}\rightarrow S$ the minimal (unramified) cover of $S$ on which a basis for the period lattice can be globally defined and we set $S_{\sigma}\rightarrow S^{*}$ to be the minimal cover of $S^{*}$ where one can define the logarithm of $\sigma$. The tower of covers is represented in the diagram:
@@
Let $\mathcal{A}\to S$ be an abelian scheme. We use Lang-Néron theorem [21, Thorem 1] to show that a non-zero algebraic section $\sigma:S\rightarrow\mathcal{A}$ which is not contained in the fixed part cannot admit a globally defined logarithm on $S(\mathbb{C})$. In fact, assume by contradiction that $\log_{\sigma}$ exists globally on $S(\mathbb{C})$, then $\frac{\sigma}{n}:=\exp\circ\left(\frac{1}{n}\log_{\sigma}\right)$ for any $n\in\mathbb{Z}\setminus\{0\}$ is an analytic section such that $n\cdot\frac{\sigma}{n}=\sigma$. This means that $\sigma$ is infinitely divisible in $\Sigma(S)$ contradicting Lang-Néron theorem.
@@
Let $\sigma:S\rightarrow\mathcal{A}$ be an analytic section and $\log_{\sigma}$ a determination of its logarithm over an open set $U\subset S(\mathbb{C})$. The section admits a globally defined logarithm on $S(\mathbb{C})$ if and only if the associated cocycle $h\mapsto(u_{1}^{h},\ldots,u_{2g}^{h})$ is a coboundary.
@@
3 Proofs of the main results
@@ 3.1 Invariance results
In this section we want to provide some preliminary results about the relative monodromy group $M_{\sigma}^{\textnormal{rel}}$ in order to prove that our main theorems are insensitive to some natural operations. First of all, let us notice that 1.1 and 1.2 are invariant under isogenies: this was proven in [33, Theorem 3.6]. In what follows we describe other invariance results.
@@
1.1 (resp. 1.2) is invariant by finite base change: in other words, if $\varphi:\widetilde{S}\rightarrow S$ is a finite morphism and $\widetilde{\mathcal{A}}\to\widetilde{S}$ is the pullback of $\mathcal{A}\rightarrow S$ via $\varphi$, then 1.1 (resp. 1.2) for $\widetilde{\mathcal{A}}\rightarrow\widetilde{S}$ implies 1.1 (resp. 1.2) for $\mathcal{A}\rightarrow S$.
@@ 3.2 Non triviality of the relative monodromy
In this section we prove 1.1. We are going to combine some topological techniques and cohomology theory in order to get some control on the relative monodromy of sections.
Let’s keep the same notation as above. Thanks to 3.1, we can assume that the abelian scheme $\phi:\mathcal{A}\to S$ is endowed with a finite surjective map $p:S\to\mathbb{A}_{g}$, where $\mathfrak{A}_{g}\to\mathbb{A}_{g}$ denotes some universal family of $g$-dimensional abelian varieties with some fixed $-\ell$-structure and without locally constant parts. Let $\sigma:S\to\mathcal{A}$ be a non-torsion section.
One of the main ideas of the proof is to construct a Kuga family associated to our abelian scheme, so that we can employ the result of Mok and To [25, Main Theorem] to obtain some constraints on the logarithm of sections. To this end, let’s consider the classifying map $p:S\to\mathbb{A}_{g}$ and look at the induced homomorphism on fundamental groups, i.e. $p_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(\mathbb{A}_{g}(\mathbb{C}),x)$ where $s\in S(\mathbb{C})$ and $x\in\mathbb{A}_{g}(\mathbb{C})$ are two base points such that $x=p(s)$. The inclusion $p_{*}\pi_{1}(S(\mathbb{C}),s)\subseteq\pi_{1}(\mathbb{A}_{g}(\mathbb{C}),x)$ gives rise to a finite unramified morphism $q^{\prime}:S^{\prime}(\mathbb{C})\to\mathbb{A}_{g}(\mathbb{C})$ that by [22, Corollary 12.19] satisfies
@@
${\mathcal{A}(\mathbb{C})}$${\mathcal{A}^{\prime}(\mathbb{C})}$${\mathfrak{A}_{g}(\mathbb{C})}$${S(\mathbb{C})}$${S^{\prime}(\mathbb{C})}$${\mathbb{A}_{g}(\mathbb{C}).}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{q^{\prime}}$ (18)
Notice that the abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ is a Kuga family by construction, since $q^{\prime}:S^{\prime}(\mathbb{C})\to\mathbb{A}_{g}(\mathbb{C})$ is an unramified morphism.
@@
We want to point out some algebraic and topological properties of the above construction. From the algebraic point of view, observe that the morphisms $q$ and $q^{\prime}$ are both finite. In fact since $p$ is finite, then $q$ and $q^{\prime}$ are quasi-finite. Since $q^{\prime}$ is a finite-sheeted (unramified) topological cover, then it is proper in the sense that the inverse image of every compact set is compact. This implies that $q^{\prime}$ is a proper morphism in the sense of scheme theory, and thus $q^{\prime}$ is finite. In particular, $q^{\prime}$ is also a separated morphism. Since $p=q^{\prime}\circ q$ is finite and $q^{\prime}$ is separated, then $q$ is proper. Since $q$ is proper and quasi-finite, then $q$ is finite.
@@
In the opposite direction, we define the trace operator
@@
as follows: we already noticed that for each $y\in S(\mathbb{C})$ the fibers $\mathcal{A}_{y}$ and $\mathcal{A}^{\prime}_{q(y)}$ are canonically identified, so for $\tau\in\Sigma^{\operatorname{an}}(S)$ we define $\operatorname{Tr}(\tau):S^{\prime}(\mathbb{C})\to\mathcal{A}^{\prime}(\mathbb{C})$ by setting
@@
where $m_{y}$ is the ramification index of $q$ at $y$. We have $\operatorname{Tr}(\Sigma(S))\subseteq\Sigma^{\prime}(S^{\prime})$; in addition, since $\mathcal{A}^{\prime}\to S^{\prime}$ is a Kuga family we have $\operatorname{Tr}(\Sigma(S))\subseteq\Sigma^{\prime}(S^{\prime})_{\textrm{tor}}$ by 2.1, where $\Sigma^{\prime}(S^{\prime})_{\textrm{tor}}$ is the group of torsion sections. Moreover, we point out that the following property holds for each holomorphic section $\tau^{\prime}:S^{\prime}(\mathbb{C})\rightarrow\mathcal{A}^{\prime}(\mathbb{C})$:
@@
where $q_{C},q^{\prime}_{C^{\prime}}$ and $p_{C}$ denote the restrictions of the corresponding maps. In this situation, assuming for the moment that $C^{\prime}$ and $D$ are still of dimension $1$, we can clearly consider the analogous notions of trace operator and pullback operator as well as the restriction of a section $\sigma:S\to\mathcal{A}$ to $C$, which will be denoted by $\sigma_{C}$. We also have analogous notions of monodromy representations and monodromy groups, e.g. we will write $\rho_{C},\theta_{\sigma_{C}},\textrm{Mon}(\mathcal{A}|_{C})$ and $M^{\textrm{rel}}_{\sigma_{C}}$ for the obvious objects associated to the scheme $\mathcal{A}|_{C}\to C$ and to the section $\sigma_{C}:C\to\mathcal{A}|_{C}$; we do the same for $C^{\prime}$. Moreover, if we denote by $i:C\hookrightarrow S(\mathbb{C})$ and $i^{\prime}:C^{\prime}\hookrightarrow S^{\prime}(\mathbb{C})$ the natural embeddings and choose base points $s\in C$ and $s^{\prime}\in C^{\prime}$, from a topological point of view we obtain the following diagram involving the induced homomorphisms between fundamental groups (for the surjectivity of the upper horizontal arrow see 3.3):
@@ 1. $(i)$
    the morphism $q_{C}:C\to C^{\prime}$ appearing in Equation 22 is finite, the curves $C$ and $C^{\prime}$ are affine and the section $\sigma_{C}:C\to\mathcal{A}|_{C}$ is non-torsion and non-isoconstant;
@@ 7. $(vii)$
    In order to construct the curve $C$ let us consider the general hyperplane section $Y=H\cap S(\mathbb{C})$ and choose a base point $s\in Y$. By Lefschetz theorem on quasi-projective varieties [18, 2.2], the natural map $\pi_{1}(Y,s)\to\pi_{1}(S(\mathbb{C}),s)$ between fundamental groups is surjective. After iterating the process of taking such generic hyperplane sections we obtain an irreducible curve $C_{1}$ with the property that the natural map $\pi_{1}(C_{1},s)\to\pi_{1}(S(\mathbb{C}),s)$ is surjective for a base point $s\in C_{1}$. The restriction of $q$ to $C_{1}$ defines a morphism $q_{C_{1}}$ which is finite: in fact, since every closed embedding is a finite morphism and $q:S\to S^{\prime}$ is a finite morphism (see 3.3), then the restriction $q_{C_{1}}$ is finite. We define $C^{\prime}:=q_{C_{1}}(C_{1})$ and $C:=q^{-1}(C^{\prime})$. Notice that $C^{\prime}$ is an affine curve since $q_{C_{1}}$ is finite and surjective; moreover, $C$ could be no longer irreducible but it is still affine because $q$ is an affine morphism (since it is finite). Furthermore, since $q:S\to S^{\prime}$ is finite then it is closed. We have a closed embedding $i:C\hookrightarrow S(\mathbb{C})$, and we can conclude that the restriction $q_{C}:C\to C^{\prime}$ is a finite morphism. Up to remove a finite number of points from $C^{\prime}$ with the corresponding preimages, we can suppose that both $C$ and $C^{\prime}$ are affine curves. In addition, by the genericity of the hyperplane sections we can assume that the section $\sigma_{C}:C\to\mathcal{A}|_{C}$ is non-torsion and non-isoconstant.
@@
Remark 3.5.
In 3.4 the fact that $\sigma_{C}$ is non-torsion follows by the genericity of hyperplane sections, but notice that actually it is also a consequence of part $(iii)$ of 3.4.
@@
Let us study the exact sequence in cohomology groups induced by Equation 24. By 2.7 and by part $(iv)$ of 3.4 we can conclude that no non-zero period can be globally defined on $C$. In other words, we get $H^{0}(C,\Lambda_{C})=0$. In addition, since $C$ is a Stein space we have
@@
By 2.11, we know that the obstruction to define a global logarithm of a section is given by some Galois cohomology class. We want to explain the interplay between the relevant Galois cohomology group and the exact sequence in Equation 25. Let’s fix the notation $G:=\pi_{1}(C,s)$ for a base point $s\in C$ and denote by $\rho_{C}:G\to\textrm{Mon}(\mathcal{A}|_{C})\subseteq\textrm{Sp}_{2g}(\mathbb{Z})$ the monodromy representation associated to the abelian scheme $\mathcal{A}|_{C}\to C$; we use the notation $\overline{h}:=\rho_{C}(h)$. With the aim of proving how the Galois cohomology interacts with the sheaf cohomology in our situation, we put the following remark in order to interpret the sections of the sheaf of periods as continuous maps on the universal cover of $C$ which are well-behaved with respect to the action of $G$.
@@
Let $(\eta_{i,j})_{i,j}$ be a cocycle in $H^{1}(\mathcal{V},\Lambda_{C})$, where $\eta_{i,j}\in\Gamma(V_{i,j},\Lambda_{C})$. We refer to the diagram of 3.6 and denote by $u_{C}:\widetilde{C}\rightarrow C$ the universal cover of $C$; moreover, let us define $U_{i}=u_{C}^{-1}(V_{i})$ and $U_{i,j}=u_{C}^{-1}(V_{i,j})=U_{i}\cap U_{j}$. By considering the pull-backs $u_{C}^{*}(\eta_{i,j})$, we obtain a cocycle $w_{i,j}$ with values in $\mathbb{Z}^{2g}$ for the covering $u_{C}^{*}(\mathcal{V})=(\pi^{-1}(V_{i}))_{i\in I}$. By the above remark, the elements $w_{i,j}$ can be viewed as continuous (i.e. locally constant) functions $U_{i,j}\rightarrow\mathbb{Z}^{2g}$, satisfying
@@
similar to the relation holding for the $w_{i,j}$. By 3.6, this fact means that the $w^{\prime}_{i}$ are of the form $w^{\prime}_{i}=u_{C}^{*}(\eta_{i})$ for suitable sections $\eta_{i}\in\Gamma(V_{i},\Lambda_{C})$, so that $\eta_{i,j}=\eta_{i}-\eta_{j}$ is a coboundary.
@@
Therefore, thanks to the exact sequence Equation 25 and 3.7 we obtain a surjective group homomorphism
@@
Notice that by 3.4 also the curve $C^{\prime}$ is a Stein space. Moreover, by Equation 9 the group $\textrm{Mon}(\mathcal{A}|_{C})$ is a finite index subgroup of $\textrm{Mon}(\mathcal{A}^{\prime}|_{C^{\prime}})$; since $\textrm{Mon}(\mathcal{A})$ acts irreducibly on the lattice of periods (see 2.6), by part $(iv)$ of 3.4 we obtain that the action of $\textrm{Mon}(\mathcal{A}^{\prime}|_{C^{\prime}})$ on the corresponding lattice of periods is still irreducible. In other words we get $H^{0}(C^{\prime},\Lambda_{C^{\prime}})=0$. Therefore, the analogous results of Equation 25 and 3.7 hold if we replace the curve $C$ by the curve $C^{\prime}$. In particular, for some base point $s^{\prime}\in C^{\prime}$ we have the surjective group homomorphism
@@
Let us point out the relationship between the cohomology groups $H^{1}(\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime}),\mathbb{Z}^{2g})$ and $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$. By part $(iii)$ of 3.4 the homomorphism $i^{\prime}_{*}:\pi_{1}(C^{\prime},s^{\prime})\to\pi_{1}(S^{\prime},s^{\prime})$ is surjective and this induces a homomorphism
@@
We are ready to prove the following theorem:
Theorem 3.10.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme such that (up to a finite base change) the modular map $p:S\to\mathbb{A}_{g}$ is a finite surjective morphism. If $\sigma\in\Sigma(S)$ is a non-torsion section, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is non-trivial.
Proof.
Thanks to 2.1 and 3.2 up to multiplying $\sigma$ by a positive integer we can suppose $\operatorname{Tr}(\sigma)=0$. Let $C$ be a curve satisfying 3.4.
We work by contradiction by assuming $M_{\sigma}^{\textrm{rel}}=\{0\}$. By part $(vii)$ of 3.4 this implies $M_{\sigma_{C}}^{\textrm{rel}}=\{0\}$, or equivalently $\ker{\rho_{C}}\subseteq\ker{\theta_{\sigma_{C}}}$. Let $q_{C*}:\pi_{1}(C,s)\to\pi_{1}(C^{\prime},s^{\prime})$ be the homomorphism appearing in Equation 23. By part $(v)$ of 3.4 we have $\ker{q_{C*}}\subseteq\ker{\rho_{C}}$; thus, we have $\ker{q_{C*}}\subseteq\ker{\theta_{\sigma_{C}}}$. Introducing the notation $\overline{G}:=q_{C*}\pi_{1}(C,s)$ we obtain that there exists a homomorphism $f:\overline{G}\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ such that the diagram
@@
To this end, let us construct some homomorphisms between cohomology groups. Since we are assuming $M_{\sigma}^{\textrm{rel}}=\{0\}$ and the map $q_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})$ is surjective by 3.3, with the same reasoning as in Equation 28 we obtain an injective homomorphism
@@
which to a cohomology class $[\tau]_{S}$ associates a cohomology class denoted with $[\tau]_{S^{\prime}}$. By 3.9 we have a homomorphism
@@
where $\tau\in\Sigma(S)$ is such that $r(\tau)=\tau_{C}$: notice that this map is well-defined by 3.4. In fact, if $\tau_{1},\tau_{2}\in\Sigma(S)$ are such that $\tau_{1,C}=\tau_{2,C}$, then $M_{\tau_{1}}^{\textrm{rel}}=M_{\tau_{2}}^{\textrm{rel}}$ by part $(vi)$ of 3.4. This implies $[\tau_{1}]_{S}=[\tau_{2}]_{S}$ and so $[\tau_{1}]_{C^{\prime}}=[\tau_{2}]_{C^{\prime}}$.

### v3

@@ 2.4 Relative monodromy of abelian logarithms
Let’s continue considering an abelian scheme $\mathcal{A}\rightarrow S$ and a non-zero section $\sigma:S\rightarrow\mathcal{A}$. Generally neither a basis of periods nor a logarithm $\log_{\sigma}$ can be globally defined on the whole of $S(\mathbb{C})$ (see 2.5 and 2.8 below), but obviously they can be globally defined on the universal cover of $S$. Studying the related monodromy problems corresponds to finding out the minimal (unramified) cover of $S$ on which both a basis of the periods and a logarithm of the section can be defined. With this in mind, we first call $S^{*}\rightarrow S$ the minimal (unramified) cover of $S$ on which a basis for the period lattice can be globally defined and we set $S_{\sigma}\rightarrow S^{*}$ to be the minimal cover of $S^{*}$ where one can define the logarithm of $\sigma$. The tower of covers is represented in the diagram:
@@
Let $\mathcal{A}\to S$ be an abelian scheme. We use Lang-Néron theorem [23, Thorem 1] to show that a non-zero algebraic section $\sigma:S\rightarrow\mathcal{A}$ which is not contained in the fixed part cannot admit a globally defined logarithm on $S(\mathbb{C})$. In fact, assume by contradiction that $\log_{\sigma}$ exists globally on $S(\mathbb{C})$, then $\frac{\sigma}{n}:=\exp\circ\left(\frac{1}{n}\log_{\sigma}\right)$ for any $n\in\mathbb{Z}\setminus\{0\}$ is a holomorphic section such that $n\cdot\frac{\sigma}{n}=\sigma$. This means that $\sigma$ is infinitely divisible in $\Sigma(S)$ contradicting Lang-Néron theorem.
@@
Let $\sigma:S\rightarrow\mathcal{A}$ be a holomorphic section and $\log_{\sigma}$ a determination of its logarithm over an open set $U\subset S(\mathbb{C})$. The section admits a globally defined logarithm on $S(\mathbb{C})$ if and only if the associated cocycle $h\mapsto(u_{1}^{h},\ldots,u_{2g}^{h})$ is a coboundary.
@@
3 Proof of the main results
@@ 3.1 Invariance results
In this section we want to provide some preliminary results about the relative monodromy group $M_{\sigma}^{\textnormal{rel}}$ in order to prove that our main theorem are insensitive to some natural operations. First of all, let us notice that 1.2 is invariant under isogenies: this was proven in [33, Theorem 3.6]. In what follows we describe other invariance results.
@@
1.2 is invariant by finite base change: in other words, if $\varphi:\widetilde{S}\rightarrow S$ is a finite morphism and $\widetilde{\mathcal{A}}\to\widetilde{S}$ is the pullback of $\mathcal{A}\rightarrow S$ via $\varphi$, then 1.2 for $\widetilde{\mathcal{A}}\rightarrow\widetilde{S}$ implies 1.2 for $\mathcal{A}\rightarrow S$.
@@ 3.2 Non triviality of the relative monodromy
In this section we prove 1.2, i.e. that $M_{\sigma}^{\textnormal{rel}}$ is non-trivial.
We are going to combine some topological techniques and cohomology theory in order to get some control on the relative monodromy of sections. Let’s keep the same notation as above and denote by $T:=p(S)\subseteq\mathbb{A}_{g}$ the image of the modular map, where $\mathfrak{A}_{g}\to\mathbb{A}_{g}$ denotes some universal family of $g$-dimensional abelian varieties with some fixed $-\ell$-structure and without locally constant parts. Thanks to 3.1, we can assume that the abelian scheme $\phi:\mathcal{A}\to S$ is endowed with a generically finite modular map $p:S\to T$. In order to simplify the notation after eventually restricting to Zariski open dense subsets we can assume that $p:S\to T$ is finite. Let $\sigma:S\to\mathcal{A}$ be a ramified section.
Remark 3.3.
Notice that the notion of ramified section is not affected by the restriction of the base. In fact the ramification property must be checked after extending the section to a maximal base. One uses the fact that for any open dense subset $U\subseteq S$ that induces an abelian scheme $\mathcal{A}|_{U}\to U$ we have $\Sigma(S)\cong\Sigma(U)$.
One of the main ideas of the proof is to construct a certain “unramified family” (i.e. unramified base change of the universal family) associated to our abelian scheme, which forces some constraints on the logarithm of ramified sections. To this end, let’s consider the classifying map $p:S\to T$ and look at the induced homomorphism on fundamental groups, i.e. $p_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(T(\mathbb{C}),x)$ where $s\in S(\mathbb{C})$ and $x\in T(\mathbb{C})$ are two base points such that $x=p(s)$. The inclusion $p_{*}\pi_{1}(S(\mathbb{C}),s)\subseteq\pi_{1}(T(\mathbb{C}),x)$ gives rise to a finite unramified morphism $q^{\prime}:S^{\prime}(\mathbb{C})\to T(\mathbb{C})$ that by [24, Corollary 12.19] satisfies
@@
${\mathcal{A}(\mathbb{C})}$${\mathcal{A}^{\prime}(\mathbb{C})}$${\mathfrak{A}_{g|_{T}}(\mathbb{C})}$${S(\mathbb{C})}$${S^{\prime}(\mathbb{C})}$${T(\mathbb{C}).}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{q^{\prime}}$ (19)
Notice that the abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ is an unramified family by construction, since $q^{\prime}:S^{\prime}(\mathbb{C})\to T(\mathbb{C})\subseteq\mathbb{A}_{g}(\mathbb{C})$ is an unramified morphism.
@@
We want to point out some algebraic and topological properties of the above construction. First we show that $q$ and $q^{\prime}$ are both finite morphisms of schemes: since $p$ is finite, then $q$ and $q^{\prime}$ are quasi-finite. Since $q^{\prime}$ is a finite-sheeted (unramified) topological cover, then it is proper in the sense that the inverse image of every compact set is compact. This implies that $q^{\prime}$ is a proper morphism in the sense of scheme theory, and thus $q^{\prime}$ is finite. In particular, $q^{\prime}$ is also a separated morphism. Since $p=q^{\prime}\circ q$ is finite and $q^{\prime}$ is separated, then $q$ is proper. Since $q$ is proper and quasi-finite, then $q$ is finite.
@@
In the opposite direction, since $q$ is a finite morphism, we define the trace operator
@@
in the following way: we already noticed that for each $y\in S(\mathbb{C})$ the fibers $\mathcal{A}_{y}$ and $\mathcal{A}^{\prime}_{q(y)}$ are canonically identified, so for $\tau\in\Sigma^{\operatorname{an}}(S)$ we define $\operatorname{Tr}(\tau):S^{\prime}(\mathbb{C})\to\mathcal{A}^{\prime}(\mathbb{C})$ by setting
@@
where $m_{y}$ is the ramification index of $q$ at $y$. We have $\operatorname{Tr}(\Sigma(S))\subseteq\Sigma^{\prime}(S^{\prime})$. Moreover, we point out that the following property holds for each holomorphic section $\tau^{\prime}:S^{\prime}(\mathbb{C})\rightarrow\mathcal{A}^{\prime}(\mathbb{C})$:
@@
Remark 3.5.
When $q$ is generically finite (but not necessarily finite) we can anyway define a trace operator for the algebraic sections $\operatorname{Tr}:\Sigma(S)\to\Sigma^{\prime}(S^{\prime})$ restricting the previous construction to the open sets where $q$ is finite.
@@
where $q_{C},q^{\prime}_{C^{\prime}}$ and $p_{C}$ denote the restrictions of the corresponding maps. In this situation, if $C^{\prime}$ and $D$ are still of dimension $1$, we can clearly consider the analogous notions of trace operator and pullback operator as well as the restriction of a section $\sigma:S\to\mathcal{A}$ to $C$, which will be denoted by $\sigma_{C}$. We also have analogous notions of monodromy representations and monodromy groups, e.g. we will write $\rho_{C},\theta_{\sigma_{C}},\textrm{Mon}(\mathcal{A}|_{C})$ and $M^{\textrm{rel}}_{\sigma_{C}}$ for the obvious objects associated to the scheme $\mathcal{A}|_{C}\to C$ and to the section $\sigma_{C}:C\to\mathcal{A}|_{C}$; we do the same for $C^{\prime}$. Moreover, if we denote by $i:C\hookrightarrow S(\mathbb{C})$ and $i^{\prime}:C^{\prime}\hookrightarrow S^{\prime}(\mathbb{C})$ the natural embeddings and choose base points $s\in C$ and $s^{\prime}\in C^{\prime}$, from a topological point of view we obtain the following diagram involving the induced homomorphisms between fundamental groups (for the surjectivity of the upper horizontal arrow see 3.4):
@@ 1. $(i)$
    the morphism $q_{C}:C\to C^{\prime}$ appearing in Equation 23 is finite, the curves $C$ and $C^{\prime}$ are affine and the abelian schemes $\mathcal{A}_{|_{C}}\to C$, $\mathcal{A}^{\prime}_{|_{C^{\prime}}}\to C^{\prime}$ have no fixed part;
@@ 7. $(vii)$
    In order to construct the curve $C$ let us consider the general hyperplane section $Y=H\cap S(\mathbb{C})$ and choose a base point $s\in Y$. By Lefschetz theorem on quasi-projective varieties [19, 2.2], the natural map $\pi_{1}(Y,s)\to\pi_{1}(S(\mathbb{C}),s)$ between fundamental groups is surjective. After iterating the process of taking such generic hyperplane sections we obtain an irreducible affine curve $C_{1}$ with the property that the natural map $\pi_{1}(C_{1},s)\to\pi_{1}(S(\mathbb{C}),s)$ is surjective for a base point $s\in C_{1}$. The restriction of $q$ to $C_{1}$ defines a morphism $q_{C_{1}}$ which is finite: in fact, since every closed embedding is a finite morphism and $q:S\to S^{\prime}$ is a finite morphism (see 3.4), then the restriction $q_{C_{1}}$ is finite. We define $C^{\prime}:=q_{C_{1}}(C_{1})$ and $C:=q^{-1}(C^{\prime})$. Notice that $C^{\prime}$ is an affine curve since $q_{C_{1}}$ is finite and surjective; moreover, $C$ could be no longer irreducible but it is still affine because $q$ is an affine morphism (since it is finite). Furthermore, since $q:S\to S^{\prime}$ is finite then it is closed. We have a closed embedding $i:C\hookrightarrow S(\mathbb{C})$, and we can conclude that the restriction $q_{C}:C\to C^{\prime}$ is a finite morphism. The abelian scheme $\mathcal{A}_{|_{C}}\to C$ has no fixed part by genericity of the hyperplane sections, and this implies that also $\mathcal{A}^{\prime}_{|_{C^{\prime}}}\to C^{\prime}$ has no fixed part.
@@
Let us study the exact sequence in cohomology groups induced by Equation 25. By 2.5 we can conclude that no non-zero period can be globally defined on $C$. In other words, we get $H^{0}(C,\Lambda_{C})=0$. In addition, since $C$ is a Stein space we have
@@
By 2.9, we know that the obstruction to define a global logarithm of a section is given by some Galois cohomology class. We want to explain the interplay between the relevant Galois cohomology group and the exact sequence in Equation 26. Let’s fix the notation $G:=\pi_{1}(C,s)$ for a base point $s\in C$ and denote by $\rho_{C}:G\to\textrm{Mon}(\mathcal{A}|_{C})\subseteq\textrm{Sp}_{2g}(\mathbb{Z})$ the monodromy representation associated to the abelian scheme $\mathcal{A}|_{C}\to C$; we use the notation $\overline{h}:=\rho_{C}(h)$. With the aim of proving how the Galois cohomology interacts with the sheaf cohomology in our situation, we put the following remark in order to interpret the sections of the sheaf of periods as continuous maps on the universal cover of $C$ which are well-behaved with respect to the action of $G$.
@@
Let $(\eta_{i,j})_{i,j}$ be a cocycle in $H^{1}(\mathcal{V},\Lambda_{C})$, where $\eta_{i,j}\in\Gamma(V_{i,j},\Lambda_{C})$. We refer to the diagram of 3.7 and denote by $u_{C}:\widetilde{C}\rightarrow C$ the universal cover of $C$; moreover, let us define $U_{i}=u_{C}^{-1}(V_{i})$ and $U_{i,j}=u_{C}^{-1}(V_{i,j})=U_{i}\cap U_{j}$. By considering the pull-backs $u_{C}^{*}(\eta_{i,j})$, we obtain a cocycle $w_{i,j}$ with values in $\mathbb{Z}^{2g}$ for the covering $u_{C}^{*}(\mathcal{V})=(\pi^{-1}(V_{i}))_{i\in I}$. By the above remark, the elements $w_{i,j}$ can be viewed as continuous (i.e. locally constant) functions $U_{i,j}\rightarrow\mathbb{Z}^{2g}$, satisfying
@@
similar to the relation holding for the $w_{i,j}$. By 3.7, this fact means that the $w^{\prime}_{i}$ are of the form $w^{\prime}_{i}=u_{C}^{*}(\eta_{i})$ for suitable sections $\eta_{i}\in\Gamma(V_{i},\Lambda_{C})$, so that $\eta_{i,j}=\eta_{i}-\eta_{j}$ is a coboundary.
@@
Therefore, thanks to the exact sequence Equation 26 and 3.8 we obtain a surjective group homomorphism
@@
Notice that by 3.6 also the curve $C^{\prime}$ is a Stein space. Moreover, by 2.5 we still obtain $H^{0}(C^{\prime},\Lambda_{C^{\prime}})=0$. Therefore, the analogous results of Equation 26 and 3.8 hold if we replace the curve $C$ by the curve $C^{\prime}$. In particular, for some base point $s^{\prime}\in C^{\prime}$ we have the surjective group homomorphism
@@
Let us point out the relationship between the cohomology groups $H^{1}(\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime}),\mathbb{Z}^{2g})$ and $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$. By part $(iii)$ of 3.6 the homomorphism $i^{\prime}_{*}:\pi_{1}(C^{\prime},s^{\prime})\to\pi_{1}(S^{\prime},s^{\prime})$ is surjective and this induces a homomorphism
@@
We are now ready to prove the first main result of the paper:
Proof of 1.2.
Let $C$ be a curve satisfying 3.6. We work by contradiction by assuming $M_{\sigma}^{\textrm{rel}}=\{0\}$. By part $(vii)$ of 3.6 this implies $M_{\sigma_{C}}^{\textrm{rel}}=\{0\}$, or equivalently $\ker{\rho_{C}}\subseteq\ker{\theta_{\sigma_{C}}}$. Let $q_{C*}:\pi_{1}(C,s)\to\pi_{1}(C^{\prime},s^{\prime})$ be the homomorphism appearing in Equation 24. By part $(v)$ of 3.6 we have $\ker{q_{C*}}\subseteq\ker{\rho_{C}}$; thus, we have $\ker{q_{C*}}\subseteq\ker{\theta_{\sigma_{C}}}$. Introducing the notation $\overline{G}:=q_{C*}\pi_{1}(C,s)$ we obtain that there exists a homomorphism $f:\overline{G}\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ such that the diagram
@@
To this end, let us construct some homomorphisms between cohomology groups. Since we are assuming $M_{\sigma}^{\textrm{rel}}=\{0\}$ and the map $q_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})$ is surjective by 3.4, with the same reasoning as in Equation 29 we obtain an injective homomorphism
@@
which to a cohomology class $[\tau]_{S}$ associates a cohomology class denoted with $[\tau]_{S^{\prime}}$. By 3.10 we have a homomorphism
@@
where $\tau\in\Sigma(S)$ is such that $r(\tau)=\tau_{C}$: notice that this map is well-defined by 3.6. In fact, if $\tau_{1},\tau_{2}\in\Sigma(S)$ are such that $\tau_{1,C}=\tau_{2,C}$, then $M_{\tau_{1}}^{\textrm{rel}}=M_{\tau_{2}}^{\textrm{rel}}$ by part $(vi)$ of 3.6. This implies $[\tau_{1}]_{S}=[\tau_{2}]_{S}$ and so $[\tau_{1}]_{C^{\prime}}=[\tau_{2}]_{C^{\prime}}$.